OpenStudy (kymber):

An object is launched at 19.6 m/s from a height of 58.8 m. The equation for the height (h) in terms of time (t) is given by h(t) = -4.9t² +19.6t + 58.8. What is the object's maximum height?

Parth (parthkohli):

Convert it to the vertex form (have you learned that yet?).

OpenStudy (anonymous):

Try to find the roots using the quadratic formula, then take the value of t half-way in between the roots and plug in. I think that should work

Parth (parthkohli):

Basically, the parabola opens downwards. So see here:\[-4.9t^2 + 19.6t +58.8 = -4.9t(t - 4) + 58.8\]

Parth (parthkohli):

Hmm actually\[-4.9t^2 + 19.6t + 58.8 = -4.9(t^2 - 4t) + 58.8\]\[= -4.9(t^2 - 4t + 4 - 4) + 58.8 = -4.9(t - 2)^2 + 54.8\]

OpenStudy (kymber):

@ParthKohli, I don't understand this at all! :( What does that mean? How did you get it? Am I supposed to solve for t?